OpenStudy (anonymous):

Given a wavelength of 121,270 millimeters, what is the wavelength in kilometers? 0.121270 12.1270 1,212.70 12,127.0

OpenStudy (anonymous):

1km = 10^6 mm Thereforeit is equal to121,270 x 10^ (-6) km = 0.12127 km

OpenStudy (anonymous):

1km = 10^6 mm Thereforeit is equal to121,270 x 10^ (-6) km = 0.12127 km

OpenStudy (chriss):

1m=1000mm 121270/1000=121.27m 1km=1000m 121.27/1000=0.12127km